OpenStudy (anonymous):

Someone please explain to me exponential growth and exponential decay? when to use it?

OpenStudy (anonymous):

Occurs when something increases or decreases as a fraction of itself, like ten percent per year: dN/dt = 0.1 N dN/N = 0.1 dt integrate log N = 0.1 t + constant constant = No, the value when t=0 log(N/No) = 0.1 t (N/No) = exp(0.1 t) for 0.1/year and t = 10 years N/No = 2.72 Instead of doubling in ten years at 10% per year, it increased 2.72x Exponential decay starts same way, but with dN/dt =- k N N/No = exp(-kt) is the result

OpenStudy (anonymous):

examples: population growth, compound interest radioactive decay